# Dark Matter from quasi-de Sitter Horizons

### v1

@@ I Introduction
Gibbons and Hawking’s foundational work Gibbons and Hawking (1977a) established and clarified the thermal nature of de Sitter horizon radiation and its thermodynamic properties, which include an entropy proportional to the horizon’s area. Subsequent studies have confirmed and extended these early results and insights, emphasizing, in particular, the importance of observer-dependent interpretations of vacuum states Lapedes (1978a). For example, Bogoliubov transformations between in-vacuum and out-vacuum states consistently predict a thermal particle spectrum for static observers in de Sitter spacetime Gibbons and Hawking (1977a); Lapedes (1978a); Kim (2012); Mottola (1985a). These results have been verified across a variety of settings, including applied quantum field theoretic treatments of massless and massive scalar fields Jiang (2020), spinor fields Jiang (2020), and semiclassical WKB tunneling calculations Medved (2002a); Parikh (2002a).
Modern approaches further extend this framework by incorporating dynamical effects and back-reaction. Studies have explored how particle production alters the background spacetime geometry, leading to a slow decay of the cosmological constant or modifications of the horizon Sekiwa (2008a); Anderson et al. (2017); Anderson and Mottola (2013a). In particular, semiclassical tunneling mechanisms have been refined to account for these phenomena, showing that particle production in de Sitter spacetime slightly deviates from perfect thermality due to horizon dynamics Zhou and Zhang (2008); Medved (2002a). Additionally, alternative approaches, including Gaussian wave packets Kim (2012), phase-integral approximations Kim (2010), and super-adiabatic particle number formalism Dabrowski and Dunne (2014), provide corroborating insights into the way radiation emerges through quantum fluctuations near horizons.
@@ II Observers in de Sitter and quasi-de Sitter spacetime
The possibility to classify ourselves as static observers within modern cosmological contexts is not straightforward. In observational cosmology, i.e. at late times, the expansion of the universe is better described by the FLRW metric, which models a quasi-de Sitter spacetime for scenarios such as inflation or the present dark-energy-dominated epoch. For FLRW comoving observers, for instance, evolving field modes and particle creation during inflation are tied to horizon crossing rather than a purely static horizon structure Ford (1987); Mottola (1985b). Examples include the production of nearly scale-invariant fluctuations in the spectrum of perturbations, as observed in inflationary cosmologies Mottola (1985b); Ford (1987). Comoving observers generalize particle creation as an evolving process tied to changes in the vacuum state, rather than as a steady-state thermal phenomenon. In addition, in cases involving vacuum decay or backreaction, deviations from the static observer perspective may arise Mottola (1985b), Anderson and Mottola (2013b).
Current cosmological observations present mixed evidence regarding whether Earth’s frame aligns with the Hubble flow. While the CMB dipole ($\sim 370$ km/s) nominally supports comoving coordinates through its kinematic interpretation, a $4.9\sigma$ tension persists between the CMB dipole direction and that of distant quasars Secrest et al. (2021). Similarly, anisotropic Hubble parameter measurements reveal a $3.9\sigma$ dipole aligned with the CMB frame Colin et al. (2019), challenging strict adherence to the cosmological principle. These anomalies coexist with otherwise consistent geometric tests: baryon acoustic oscillation measurements confirm isotropy to within $2\%$ Alam et al. (2021), and gravitational redshift surveys align with $\Lambda$CDM predictions Wojtak et al. (2011).
Future observations could resolve this ambiguity through novel kinematic probes. Redshift drift measurements with next-generation radio telescopes like SKA could detect peculiar velocity signatures at the $\Delta z\sim 10^{-10}$ level Darling (2021), directly constraining non-Hubble motions. Similarly, direction drift effects from high-precision astrometry (e.g., Gaia) may distinguish parallax evolution in comoving frames from aberration effects in static frames Quercellini et al. (2009). Upcoming peculiar velocity surveys (DESI, LSST) will test whether local flows align with $\Lambda$CDM predictions or indicate residual frame-dependent accelerations Mandelbaum et al. (2018).
While current data broadly supports the comoving paradigm, the $\sim 600$ km/s “dark flow” component Kashlinsky et al. (2008) and Hubble tension ($\sim 5\sigma$) suggest either unmodeled systematics or fundamental frame effects. Definitive resolution awaits measurements of redshift drift anisotropies ($\sim 10\%$ at $z\sim 0.1$) and CMB spectral distortions that could disentangle kinematic versus intrinsic dipole origins Aghanim et al. (2016).
@@
We also note that the transition from inflation to a second accelerated expansion phase in the very early universe can be achieved through several possible mechanisms: The inflaton field $\phi$ could evolve to a different region of its potential $V(\phi)$, where $-1<w<-1/3$ Linde (1982); A separate field $\psi$ with a suitable potential $U(\psi)$ could become dominant, driving a late accelerated expansion Kofman et al. (1994); finally, modifications to Einstein’s field equations at high energies could induce accelerated expansion without requiring an additional field, see e.g. Starobinsky (1980). Each of these mechanisms could potentially lead to a phase of accelerated expansion with $-1<w<-1/3$ immediately following inflation, at temperatures $T\gg 10$ MeV, before the onset of radiation domination.
@@ III Particle production from quasi-de Sitter horizons
Figure 1: The energy density of radiation (blue, right) and of fluids with three different equations of state parameters $w$ and transition-to-radiation temperatures $T_{r}$.
@@
with $g_{*}=106.75$ and $\bar{M}_{\rm Pl}\simeq 2.4\times 10^{18}$ GeV the reduced Planck mass. We show for clarity the assumed energy density structure, for a few choices of $w$ and $T_{r}$ in Fig. 1.

### v2

@@ I Introduction
Gibbons and Hawking’s foundational work Gibbons and Hawking (1977a) established and clarified the thermal nature of de Sitter horizon radiation and its thermodynamic properties, which include an entropy proportional to the horizon’s area. Subsequent studies have confirmed and extended these early results and insights, emphasizing, in particular, the importance of observer-dependent interpretations of vacuum states Lapedes (1978a). For example, Bogoliubov transformations between in-vacuum and out-vacuum states consistently predict a thermal particle spectrum for static observers in de Sitter spacetime Gibbons and Hawking (1977a); Lapedes (1978a); Kim (2012); Mottola (1985). These results have been verified across a variety of settings, including applied quantum field theoretic treatments of massless and massive scalar fields Jiang (2020), spinor fields Jiang (2020), and semiclassical WKB tunneling calculations Medved (2002a); Parikh (2002a).
Modern approaches further extend this framework by incorporating dynamical effects and back-reaction. Studies have explored how particle production alters the background spacetime geometry, leading to a slow decay of the cosmological constant or modifications of the horizon Sekiwa (2008a); Anderson et al. (2017); Anderson and Mottola (2013). In particular, semiclassical tunneling mechanisms have been refined to account for these phenomena, showing that particle production in de Sitter spacetime slightly deviates from perfect thermality due to horizon dynamics Zhou and Zhang (2008); Medved (2002a). Additionally, alternative approaches, including Gaussian wave packets Kim (2012), phase-integral approximations Kim (2010), and super-adiabatic particle number formalism Dabrowski and Dunne (2014), provide corroborating insights into the way radiation emerges through quantum fluctuations near horizons.
@@ II Observers in de Sitter and quasi-de Sitter spacetime
The particle production mechanism proposed in this work, and detailed upon in the next section, shares conceptual ancestry with quantum field theoretic treatments of particle creation in time-dependent backgrounds, as developed in the seminal works of Parker Parker (1968, 1969), Fulling Fulling (1973), Hawking Hawking (1974, 1975), and Gibbons and Hawking Gibbons and Hawking (1977c). In those approaches, field modes are globally quantized with respect to vacua defined at asymptotic past and future times. Particle production arises due to nontrivial Bogoliubov transformations between these in/out vacua, leading to a squeezed state interpretation of the evolving field. While this framework is well-established and rigorous, it typically requires knowledge of global spacetime structure and often focuses on idealized backgrounds such as exact de Sitter space or power-law inflation.
Our approach is complementary in spirit but operationally distinct. Rather than working with global vacuum modes, we consider local particle production associated with comoving modes as they cross the Hubble horizon. In this picture, particles are “created” when the physical wavelength of a comoving mode becomes comparable to the Hubble radius, and their phase-space distribution is determined by the scale-dependent redshifting governed by the evolving Hubble parameter $H(t)$. This yields a differential production rate that can be integrated over cosmological time, capturing contributions from multiple epochs.
A key difference lies in the physical interpretation of the produced particles. In standard QFT treatments, the particle concept is intrinsically observer-dependent and often tied to an adiabatic vacuum definition, leading to ambiguities in non-stationary backgrounds. In contrast, our framework assumes a well-defined particle species with mass $m$ and negligible interactions, allowing us to track their comoving number density and kinetic energy after horizon exit. This makes our approach particularly suited for modeling cold dark matter production, where the focus is not on particle number ambiguity, but on the emergence of a relic population with well-defined momentum and abundance. Our formalism naturally encodes redshifting effects and allows for continuous tracking of particle phase-space densities post-production. This facilitates a more direct computation of dark matter observables such as velocity dispersion and clustering properties, which are not easily extracted from standard QFT treatments. While both approaches aim to quantify particle creation in curved spacetimes, our horizon-based formulation emphasizes local, semiclassical, and cosmologically motivated features that are particularly relevant for scenarios involving cold, nonthermal relics such as dark matter.
In discussing horizon-scale particle production, it is important to specify the class of observers with respect to which particle content is defined. While particle production in standard FRW cosmology is typically described from the point of view of comoving geodesic observers, effects analogous to the Gibbons–Hawking mechanism in de Sitter space suggest that a static observer—one associated with the apparent horizon—may provide a more natural frame for identifying a locally thermal particle flux. In our framework, the production of dark matter occurs well before the formation of local structures such as galaxies or the Earth, but the observer-dependence of particle content remains conceptually relevant. Our assumption is that the particles produced near the horizon can be meaningfully counted from the viewpoint of an effective observer anchored to the evolving causal structure, enabling a semiclassical description of their phase-space distribution that persists to late times.
@@
We also note that the transition from inflation to a second accelerated expansion phase in the very early universe can be achieved through several possible mechanisms: The inflaton field $\phi$ could evolve to a different region of its potential $V(\phi)$, where $-1<w<-1/3$ Linde (1982); A separate field $\psi$ with a suitable potential $U(\psi)$ could become dominant, driving a late accelerated expansion Kofman et al. (1994); another example is the model discussed in Ref. Fichet et al. (2024), a holographic setup which can be thought either as a model of non-conformal strongly-interacting sector or as a non-AdS braneworld model. In either viewpoint, the model predicts the existence of an extra fluid in the 4D Friedmann equation with equation of state $-1<w<-1/3$; finally, modifications to Einstein’s field equations at high energies could induce accelerated expansion without requiring an additional field, see e.g. Starobinsky (1980). Each of these mechanisms could potentially lead to a phase of accelerated expansion with $-1<w<-1/3$ immediately following inflation, at temperatures $T\gg 10$ MeV, before the onset of radiation domination.
Here, both for simplicity and seeking to be model-independent, we just assume that the quasi-de Sitter phase with equation of state $-1<w<-1/3$ ends in a standard reheating transition, during which the dominant vacuum-like energy density is converted into a thermal radiation bath. This process, while not modeled explicitly here, involves the decay or dissipation of the background component into relativistic particles and is accompanied by entropy production. We adopt the instantaneous reheating approximation, in which the energy transfer occurs rapidly at a characteristic temperature $T_{r}$, marking the onset of radiation domination. The entropy density $s(T_{r})$ used to normalize the final dark matter yield reflects the entropy produced during this transition.
@@ III Particle production from quasi-de Sitter horizons
Figure 1: Cumulative contribution to the universe’s energy density of radiation (blue, right) and of fluids with three different equations of state parameters $w$ as a function of temperature. Note: for $T>T_{r}$, the temperature axis should not be interpreted as a physical thermal bath temperature, but rather as a proxy for the inverse scale factor, $T\propto 1/a$, calibrated such that $T=T_{r}$ marks the onset of radiation domination.
@@
with $g_{*}=106.75$ and $\bar{M}_{\rm Pl}\simeq 2.4\times 10^{18}$ GeV the reduced Planck mass. We show for clarity the assumed energy density structure, for a few choices of $w$ and $T_{r}$ in Fig. 1. Note that for $T>T_{r}$, the temperature axis should not be interpreted as a physical thermal bath temperature, but rather as a proxy for the inverse scale factor, $T\propto 1/a$, calibrated such that $T=T_{r}$ marks the onset of radiation domination.
@@
It is important to note that during the quasi-de Sitter epoch preceding reheating, the universe is not in thermal equilibrium and the usual concept of entropy density $s=\frac{2\pi^{2}}{45}g_{*}T^{3}$ does not apply in a strict sense. The background energy density is dominated by a vacuum-like component with negligible radiation. Therefore, temperatures $T>T_{r}$ appearing in the integrand of Eq. (6) should not be interpreted as physical radiation temperatures, but rather as redshift markers tied to comoving scales via $T\propto 1/a$, normalized to the post-reheating radiation bath.
We use the entropy density $s$ only at or below the reheating temperature $T_{r}$ to compute the final yield. In this context, $s(T_{r})$ reflects the standard radiation-dominated value after reheating is complete. The appearance of $T$ in the integral is thus a bookkeeping device, not an assertion of thermal equilibrium prior to reheating [^4] This approach is consistent with other analyses in which particle production occurs before reheating, such as gravitational production of dark matter in inflationary cosmologies (e.g., Chung et al. (1998); Garny et al. (2016))..